OpenStudy (anonymous):

help solve for x 3x2 + 2x = 7

OpenStudy (cwrw238):

3x^2 + 2x - 7 = 0 use quadratic formula

OpenStudy (anonymous):

3x^2+2x=7 first solve 3x^2 3x^2= 9x so now the equation becomes 9x+2x=7 add 9x+2x equation becomes 11x=7 divide 11x by 11 to get the x by itself, do the same to 7 (divide by 11) your answer is x=7/11

OpenStudy (anonymous):

thanks guys!!!!!

OpenStudy (anonymous):

you are welcome! :)

OpenStudy (cwrw238):

??????????? how do u get 3x^2 = 9x??

OpenStudy (anonymous):

huh just now looked at that

OpenStudy (cwrw238):

3^2 *x = 9x NOT 3x^2

OpenStudy (cwrw238):

you need to use quadratic formula x = [-b +- sqrt(b^2 - 4ac)] / 2a

OpenStudy (cwrw238):

here a = 3, b = 2 and c = -7